OpenStudy (anonymous):

howdo you evaluate the expression if the number inthe square root has an exponent?

OpenStudy (anonymous):

divide that exponent by 2 and you are all set

OpenStudy (anonymous):

\[\sqrt{x^{16}}= x^8\]

OpenStudy (anonymous):

thanks (:

OpenStudy (anonymous):

what if it has an exponent 3 befor the equation? haha sorry i dont know the right words

OpenStudy (anonymous):

\[\sqrt{x^3}\] like this?

OpenStudy (anonymous):

no.its right beforethe squareroot

OpenStudy (anonymous):

\[3\sqrt{x}\]?

OpenStudy (anonymous):

yes (:

OpenStudy (anonymous):

Then it is as simplified as it can be.

OpenStudy (anonymous):

so i just leave it like that?

OpenStudy (anonymous):

Yes, Unless your problem specifically ask for something else

OpenStudy (anonymous):

thankyou somuch! (:

OpenStudy (anonymous):

You're welcome

OpenStudy (anonymous):

okay.onemorequestion. the equation says ^3squareroot(n^75). so does the 3 meani just leave it alone or divide by 2?

OpenStudy (anonymous):

\[3\sqrt{n^{75}}\]?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Try using the "Equation" button under the box where you are typing. However, if I understood correctly, you have: \[\sqrt[3]{n^{75}}\]

OpenStudy (anonymous):

\[3n^{75/2}\]

OpenStudy (anonymous):

ooh gosh, i feel dumb now. haha im new to this placde... but yeah..and thanks

OpenStudy (anonymous):

If it is a cube root \[\sqrt[3]{n^{75}}\] you would divide the exponent by 3 n^(75/3)

OpenStudy (anonymous):

For future reference a square root is the same as raising to the (1/2) power A cubed root is raising to the (1/3) power, etc.

OpenStudy (anonymous):

thankyou guys so much! (:

OpenStudy (anonymous):

For instance: \[9^{0.5}=\sqrt9=3\]